Company A and B

Statement of ComprehensiveIncome

For theyear ended, December 31,2020

-(netsales(Company A) - 1,703,960.50

-netsales(Comapny B) - 372,941.80

-cost of sales(Company A) - 1,188,539.00

-cost of sales(Company B) - 265,100.00

-sellingExpenses(Company A) - 332,502.50.

-sellingExpenses(Company B) - 86,900.00

-General&AdministrativeExpenses(Company A) - 59,950.00.

-General&AdministrativeExpenses(Company B 7,315.00

-InterestExpenses(Company A) - 9,878.00.

-InterestExpenses(Company B) - 2,477.20

-NetIncome(Company A) - 113,091.00.

-NetIncome(Company B) - 11,149.60

Answer the following:

1. Compute for the profitability ratios of Company A and B.

2. Compute for operational ratios of company A and B.

3. Compute for the financialhealth ratios of CompanyA and B.


1
Expert's answer
2021-10-21T10:48:52-0400

1] Net profit Margin A = Net profit/Total sales

"A = 113,091\/1703960.50 *100 = 6.64" %

"B= 11,149.60\/372,941.80 = 3%" %

2] Operating ratio = operating expenses/net sales

"A = (332502.50+59950)\/1703960.50"

"A= 392,452.50\/1,703,960.50*100"

"A= 23.03%" %

"B= (86,900+7,315)\/372,941.80"

"B=94,215\/372,941.80*100"

"B= 25.26%" %

3] Debt to Income ratio = Debt / Net income *100

"A= 9,878\/113,091*100"

"A=8.74%" %

"B=2,477.20\/11,149.60*100"

"B= 22.22%" %